OpenStudy (anonymous):

Find the slope of the line passing through (4, 9) and (-2, 9) A. 0 or B. undefined

OpenStudy (solomonzelman):

\[\huge {\color{blue} {m=\frac{y_1-y_2}{x_1-x_2}} }\] plug in your points.

OpenStudy (anonymous):

I don't understand how to plug in. I was absent and the teacher gave me this as home work which I don't understand. :/

OpenStudy (solomonzelman):

OK. \[\huge {\color{blue} {(x,y)} }\]this is a REGULAR COORDINATE. plug in the \[\huge {\color{blue} {4} }\huge {\color{blue} {\And-2} }\]for X s and \[\huge {\color{blue} {9\And9} }\]for Y s.

OpenStudy (anonymous):

Ok brb

OpenStudy (solomonzelman):

??? I'll give you an example, ready?

OpenStudy (anonymous):

Sure

OpenStudy (solomonzelman):

Find the slope of the line passing through (3, 4) and (-1, 5).

OpenStudy (solomonzelman):

\[\huge {\color{red} {m=\frac{y_1-y_2}{x_1-x_2}} }\]this is the slope formula. your points are, (3, 4) and (-1, 5) plug them in LIKE THIS.

OpenStudy (solomonzelman):

\[\huge {\color{blue} {\frac{4-5}{3-(-1)}⇒\frac{4-5}{3+1}⇒\frac{-1}{4}⇒-\frac{1}{4}} }\]

OpenStudy (solomonzelman):

So the slope is, \[\huge {\color{green} {-\frac{1}{4}} }\]

OpenStudy (solomonzelman):

Do your question the same way.